- Press Release
- Dec 7, 2022
NASA Space Station On-Orbit Status 7 February 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Underway: Week 16 of Increment 10.
FE Salizhan Sharipov conducted the fourth of the current science sessions with the Russian/German TEKh-20 Plasma Crystal-3 (PK-3) payload, modified from last Friday’s run. [Right after wakeup, Salizhan again activated the PK-3/N turbopump in the Service Module (SM) Transfer Compartment (PkhO), to keep the work chamber (ZB) in the SM RO (Work Compartment) at a vacuum. Then he set up the experiment, supported by tagup with the ground via S-band. Video recording began ~5 min after experiment initiation. The experiment, running in automatic mode, is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective today was to obtain highly ordered (crystal) plasma dust at various pressures, particle quantities and disturbances of varying degree. The experiment used large particles as well as a mixture of large and small particles. The turbopump will be deactivated tonight at ~4:25pm EST.]
Later, Sharipov unstowed and set up the equipment for the BTC-10 “Cardiocog” experiment, then performed his third session of the procedure. [Originally part of Pedro Duque’s VC5 “Cervantes” science program, CARDIOCOG-2, involving new equipment that arrived with Yuri Shargin in October ’04, studies changes in the human cardiovascular system in micro-G, expressed in the peripheral arteries, and the vegetative regulation of arterial blood pressure and heart rate. For the experiment, Salizhan had to take systolic & diastolic blood pressure measurements and heart rate data manually, using the Tensoplus sphygmomanometer and storing the data on the French EGE-2 laptop. The experiment also includes a 5-minute cognitive stress test with a numbers table, with the results called out for recording.]
Continuing long-term outfitting activities on the Russian ASN-M satellite navigation system in the SM, planned for the entire week until 2/11, the FE today uninstalled and removed the ASN’s backup antenna hardware (ASN-2401-2), then searched for and installed four newly delivered receiver module units (NPM-1,2,3,4). The work was supported by S-band tagup with ground specialists. [In tomorrow’s ASN activity, Salizhan will route and connect the onboard cable network (BKS) to the NPM units, followed on 2/9 by installing two navigation computing units (NVM-1,2). On 2/10, Salizhan will then connect the ASN-M to the SUBA onboard equipment control system and on 2/11 to the SBI onboard measuring system), completing the installation. Although not yet operational, ASN-M uses GLONASS satellites (the Russian GPS equivalent) to update the ISS state vector (SV, position & velocity plus time) without using the ground (which currently has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time. The ASN equipment was originally factory-installed in the SM but was found faulty and had to be returned to the ground. After repair it was shipped again to the station on Progress 11 and re-installed by Yuri Malenchenko on 7/8/03, followed by various troubleshooting attempts to date.]
On the Human Research Facility (HRF) rack, CDR/SO Chiao conducted another checkout of the MedOps cardiac defibrillator. [This periodic routine task is scheduled as soon as possible from Expedition start and every 60 days thereafter. For the checkout, the defib is connected to the 120V outlet, equipped with its battery (#1010) and then allowed to charge, for about five seconds, to a preset energy level (e.g., 100 joules). After the button-triggered discharge, a console indicator signals success or failure of the test. The pacing signal is downlinked via S-band for 2 min. The HRF is powered down afterwards.]
Leroy had an hour reserved for consolidating flight crew equipment (FCE) contents of two half-CTBs (crew transfer bags) to one CTB, marked for return to Houston on LF-1 (STS-114).
The CDR also transferred used alkaline batteries to five CDMK (carbon dioxide monitoring kit) and CSA-CP (compound specific analyzer-combustion products) discharged batteries packs.
Following ground instructions, Salizhan took length measurements of the power cable for the A-2 air-scrubbing filter, pulled from FGB stowage and subsequently restowed. The A-2 box with folded cable was also photographed for downlink. [The A-2 will be used in the European ATV (Automated Transfer Vehicle), and today’s activity was in preparation for ATV operations.]
Leroy performed the periodic on-orbit load calibration of the SchRED (Schwinn resistive exercise device) Flexpack canisters #1002 & #1004 that came up on Progress 15. (Done last 12/29/04). [The calibration of the cans re-establishes the relationship of specific load settings with a specific number of pulls per setting, followed by recording of the load values measured with a calibration tool and steel handles from the on-orbit calibration kit.]
Working off the voluntary Russian task list, the FE conducted the regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment that researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.
Chiao completed his daily check of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC’s temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]
The CDR conducted the routine SOZh/ECLSS servicing/inspection in the SM, including ASU toilet facilities maintenance and the weekly checkup on the BRPK air/liquid condensate separator apparatus. He also prepared the regular IMS “delta” file for the daily automated export/import to the three IMS databases on the ground.
The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 4 of a new set) and one hour on VELO.
Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
Shuttle prepack and clean-up activities in the U.S. segment by Leroy today focused on the Lab’s portside Zone 4 rack (LAB1P4), a zero-G storage rack (ZSR) that is slated to return on LF-1.
The Elektron O2 generator failed twice over the weekend and was re-activated both times in 50-amp mode. The machine was deactivated today to stay below the flight rule upper limit of ppO2 = 24.1% . It will remain off for the rest of the month per the plan. [This will accommodate regeneration of the BMP micropurification unit (which shares its vacuum vent line with Elektron). TsUP/Moscow is planning a 10-mmHg repress on either 2/14 or 2/15, and a 15-mm Hg repress on 2/25 from Progress 16P storage.]
The CDR also assisted upcoming ground-commanded TCS (thermal control system) activities for tomorrow’s Metox (metal oxide) regeneration by connecting a jumper from the ITCS LTL (internal TCS low temperature loop). Later, MCC-H changed the LTL temperature setpoint to 9.4 degC.
The ground also turned on radiator truss heaters and the ETCS (external TCS) ammonia PCVP (pump control valve package) for a power-on test.
Preparatory to the upcoming Metox regeneration, today’s Elektron deactivation and the planned cabin air refreshes with Progress O2, the MCA (major constituents analyzer) was taken by the ground through a 30-min. “zero” calibration. MCC-H is using MCA measurements to closely monitor cabin ppO2 (oxygen partial pressure) and other components.
Also in preparation for tomorrow’s high-temp bakeout of the two spent Metox canisters in the U.S. Airlock (A/L), MCC-H will turn on the CDRA (carbon dioxide removal assembly) at ~4:00pm EST, then activate the A/L CCAA (common cabin air assembly) air conditioner around midnight, with temperature setpoint lowered to 7.7 degC.
MCC-H reported today that SIGI-1 (Space Integrated GPS/Inertial Navigation System #1) is now fully operational. Also, ground specialists are developing and fine-tuning procedures for the removal of the failed heat exchanger (HX) in the A/L by Chiao and Sharipov. [A new, improved HX will be delivered on Progress 17.]
No CEO (crew earth observation) targets today.
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- ISS Reboost — 2/15 (~8:22pm EST, ~1.8 m/s; phasing for 17P launch);
- Progress M-51 (16P) undocking & destructive reentry — 2/26/05;
- Progress M-52 (17P) launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz TMA-6 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
- Soyuz TMA-5 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
- Progress M-53 (18P) launch — 6/10/05;
- Progress M-54 (19P) launch — 8/24/05;
- Soyuz TMA-7 (11S) launch — 9/27/05.
ISS Orbit (as of this morning, 7:08am EST [= epoch]):
- Mean altitude — 357.1 km
- Apogee height — 363.4 km
- Perigee height — 350.8 km
- Period — 91.68 min.
- Inclination (to Equator) — 51.64 deg
- Eccentricity — 0.00093
- Solar Beta Angle — -56.7 deg (magnitude decreasing)
- Orbits per 24-hr. day — 15.70
- Mean altitude loss in last 24 hours — 60 m
- Revolutions since FGB/Zarya launch (Nov. 98) — 35529
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.